OpenStudy (pulsified333):

Suppose that you flip a coin 13 times. What is the probability that you achieve at least 5 tails?

OpenStudy (anonymous):

thats a binomial probability problem

OpenStudy (anonymous):

lol your math teacher must hate you

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

at least 5 means 5 or 6 or 7 or 8 or 9 or 10 or 11, or 12,or 13

OpenStudy (pulsified333):

No he just gives us insane amounts of homework while the other teachers at the university give like half of what he gives us

OpenStudy (anonymous):

you can compute the probability you get 0,1,2,3,4 and subtract that total from 1 it is still going to be a pain

OpenStudy (anonymous):

\[P(x=0)=\left(\frac{1}{2}\right)^{13}\]

OpenStudy (anonymous):

naw chill just use the formula of sumation

OpenStudy (anonymous):

\[P(x=1)=13\left(\frac{1}{2}\right)^{13}\]

OpenStudy (anonymous):

use you calculator to plug it into a summation

OpenStudy (pulsified333):

haven't learned that formula yet

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

\[P(x=2)=\binom{13}{2}\left(\frac{1}{2}\right)^{13}\]

OpenStudy (anonymous):

ok how about the nCr * p^r * q^(n-r) ?

OpenStudy (pulsified333):

yeah i know that

OpenStudy (anonymous):

ok now plug it in for the values 1, 2, 3, 4, 5 and add them then subtract that sum from 1

OpenStudy (anonymous):

1, 2, 3, 4, 5 being your success rate

OpenStudy (anonymous):

cool?

OpenStudy (pulsified333):

yeah

OpenStudy (pulsified333):

but what does (13 2) mean?

OpenStudy (anonymous):

13 tries and there are only 2 possible outcomes

OpenStudy (pulsified333):

oh okay

OpenStudy (anonymous):

subtract this number from 1 http://www.wolframalpha.com/input/?i=%281%2B13%2B78%2B286+%2B715%29%2F2^%2813%29

OpenStudy (anonymous):

lol

OpenStudy (pulsified333):

thanks :D